OpenStudy (anonymous):

The volume of a cylinder is 980 pi in^3. The height of the cylinder is 20 in. What is the radius of the cylinder?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

are you able to help i thought all i had left to do was what we worked on yesterday but i still have Volumes of Prisms and Cylinders Volumes of Pyramids and Cones Surface Area and Volume of Spheres Areas and Volumes of Similar Solids equations to do 10 each so basically 40 by tomorrow

Directrix (directrix):

The volume of a right circular cylinder can be found by this formula: V =B*h where B is the area of the circular base (pi*r^2) and h is the height of the cylinder.

OpenStudy (anonymous):

i need help on that

OpenStudy (anonymous):

5.38? @Directrix

Directrix (directrix):

I made a bad division. Sorry. V = pi * r^2 * h 980 pi = pi r^2 (20) 49 = r^2 @desfields901 What the square root of 49?

OpenStudy (anonymous):

7

Directrix (directrix):

Yes, so r = 7

OpenStudy (anonymous):

thats it?

Directrix (directrix):

What is the radius of the cylinder? To my thinking, that is it.

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

could you explain to me the volume thing cause i got 4 more problems but they are hard to understand

Directrix (directrix):

The volume formula changes depending on what type solid you are asked to find. First, look up the formula for the solid in the problem. Fill in the numbers you are given and find the missing one. @desfields901